# From Selection to Generation: A Survey of LLM-based Active Learning

### v1

@@ 1 Introduction
Despite the success of traditional active learning methods, the advent of Large Language Models (LLMs) with remarkable reasoning and generation capabilities creates a new paradigm of active learning. For example, as illustrated in Figure 1, instead of solely relying on a predefined metric to query data instances, such as uncertainty Kirsch et al. (2019); Diao et al. (2023) or diversity Agarwal et al. (2020); Citovsky et al. (2021), LLMs can now be used to select most informative instances after reasoning or even generate entirely new instances that are better suited for efficient model training Bayer and Reuter (2024); Parkar et al. (2024); Yang et al. (2024). Moreover, with the collected informative data instances, LLMs also enable new data annotation schemes by collaborating with a human labeler or directly simulating a human labeler Xiao et al. (2023); Kholodna et al. (2024); Wang et al. (2024), which further reduces manual annotation efforts.
However, in spite of the immense potential of LLMs for active learning—particularly in high-quality data acquisition and annotation for efficient model training—existing surveys primarily focus on traditional active learning techniques, necessitating an up-to-date review of how LLMs have advanced AL in recent years. In this paper, we address this gap by presenting the first comprehensive survey of LLM-based AL techniques, which introduces a unifying taxonomy centered on the two main components of active learning: Querying (selecting or generating unlabeled instances) and Annotation (assigning labels). Table 1 and Figure 2 provide an overview of the proposed taxonomy. Table 2 further provide an intuitive comparison of existing LLM-based AL methods from the aspects of taxonomy and applications. Guided by this taxonomy, our survey organizes and systematically reviews recent works across key aspects of LLM-based active learning as follows.
@@
  • Termination5): We discuss how recent works consider LLM costs for terminating the AL loop.
@@ 2 What is LLM-based Active Learning?
In this section, we introduce some preliminaries. We start with basic notations of active learning and then introduce the LLM-based active learning loop consisting of five main steps.
Active Learning
Let $\mathcal{U}=\{\bm{\mathrm{x}}_{i}\}_{i=1}^{N}$ be a pool of $N$ unlabeled instances, where $\bm{\mathrm{x}}_{i}\in\mathcal{X}$ are feature vectors in the input space $\mathcal{X}$. Furthermore, let =$\mathcal{L}=\{(\bm{\mathrm{x}}_{i},y_{i})\}_{i=1}^{M}$ be a labeled dataset, where $y_{i}\in\mathcal{Y}$ are the corresponding labels from the label space $\mathcal{Y}$ where the size of the labeled set $|\mathcal{L}|=M$ grows as more data instances are labeled. We are also given an annotation budget $k$ that limits the number of instances that can be labeled by the annotator. We have a predictive model $f_{\theta}:\mathcal{X}\to\mathcal{Y}$ parameterized by $\theta$ to be iteratively trained using the labeled data $\mathcal{L}$. The objective of traditional AL is to efficiently select and label a subset of the unlabeled data instances $\bm{\mathrm{x}}_{i}\in\mathcal{U}$ to maximize the performance of model $f_{\theta}$ before reaching the annotation budget $k$.
@@
LLM-based AL shares a similar goal. However, it is not bounded by the unlabeled set $\mathcal{U}$, but can also use an LLM $\mathcal{M}$ to generate entirely new data instances denoted as $\bm{\mathrm{x}}^{\prime}\not\in\mathcal{U}$ as well as generating corresponding labels $y^{\prime}$ by simulating a human annotator. We formulate in details the LLM-based AL loop as follows.
  • Initialize: For a good starting point of the active learning loop, an LLM $\mathcal{M}$ can be used to annotate an initial set of labels or generate an initial dataset $\mathcal{L}_{\text{init}}$ to warm up the target model $f_{\theta}$. This approach overcomes the cold start problem that traditional AL methods face when there is no labeled data instance available and the initial model $f_{\theta}$ does not offer sufficient information for selecting informative data instances.
  • Query: With an initialized model $f_{\theta}$, the Querying3) module is implemented to acquire the most informative data instances. Extending traditional AL methods that select from the unlabeled set $\mathcal{U}$ with certain uncertainty or diversity metrics, the LLM $\mathcal{M}$ can now be used to select instances $\bm{\mathrm{x}}_{i}\in\mathcal{U}$ either by scoring or directly choosing, augment existing examples such as generating paraphrases or contrast examples, or even generate entirely new labeled examples $\bm{\mathrm{x}}^{\prime}\not\in\mathcal{U}$.
  • Annotate: The acquired data instances $\bm{\mathrm{x}}_{i}$ or $\bm{\mathrm{x}}^{\prime}$ are then passed to the Annotation4) module to obtain corresponding labels $y_{i}$ or $y^{\prime}$. The LLM $\mathcal{M}$ can be used either in conjunction with, to augment, or even to entirely replace the human labeler. The labeled instances $\bm{\mathrm{x}}_{i}$ are then excluded from the unlabeled set $\mathcal{U}$ and added to the labeled set $\mathcal{L}$ with their labels $(\bm{\mathrm{x}}_{i},y_{i})$ or $(\bm{\mathrm{x}}^{\prime},y^{\prime})$.
  • Train: With newly annotated data instances added to the labeled dataset $\mathcal{L}$, the target model $f_{\theta}$ is trained with a step update on its parameters $\theta$ with respect to the loss $\ell((\bm{\mathrm{x}}_{i},y_{i}),{f_{\theta}})$. The updated model $f_{{\theta}}$ is then used to provide information for the querying module in the next iteration before a termination criterion is met.
  • Terminate: The active learning loop reaches Termination5) when a fixed annotation budget $k$ is reached, or when some property of the model $f_{\theta}$ being optimized, such as convergence, is satisfied. With LLMs, the budget can be not only based on the availability or cost of human annotators, but also the cost of prompting the LLM or a combination of both.
@@
5:while not Terminate$(k,f_{\theta},\mathcal{M})$ do
@@
8:     $\mathcal{U}~{}\leftarrow~{}\mathcal{U}\;\backslash\;\{\bm{\mathrm{x}}\},~{}~{}\mathcal{L}~{}\leftarrow~{}\mathcal{L}\cup\{(\bm{\mathrm{x}},y)\}$
9:     $f_{\theta}~{}\leftarrow~{}\texttt{Train}(\mathcal{L})$
10:return ${}f_{{\theta}^{*}},~{}\mathcal{L}$
@@ 3 Query: From Selection to Generation
Active learning (AL) fundamentally seeks to maximize model performance with minimal annotation cost by carefully selecting the most informative examples. Traditionally, this process has relied on uncertainty-based and diversity-based metrics Settles (2011); Wang and Shang (2014); Geifman and El-Yaniv (2017); Citovsky et al. (2021). With the advent of large language models (LLMs), however, the paradigm is evolving from merely selecting examples from a fixed unlabeled pool to also generating new, high-value queries on demand.
@@ 3.1 Traditional Selection
In this section, we briefly survey some of the key existing active learning query selection strategies. The uncertainty metric captures how uncertain the model is in predicting the label of the example, whereas the diversity metric captures how different the chosen example is from the already selected examples in the labeled pool. Uncertainty-based methods, such as Least Confidence (Settles, 2009, 2011; Wang and Shang, 2014), Margin Sampling (Tong and Koller, 2001; Balcan et al., 2009; Settles, 2009), and Max-Entropy (Wang and Shang, 2014; Kremer et al., 2014; Diao et al., 2023), quantify how unsure a model is about its predictions. Complementary to these, diversity-based strategies—like Coreset (Sener and Savarese, 2017; Geifman and El-Yaniv, 2017; Citovsky et al., 2021) and CDAL (Agarwal et al., 2020)—ensure that selected examples cover varied regions of the input space. Hybrid approaches, including Badge (Ash et al., 2019, 2021) and Bald (Kirsch et al., 2019; Gal et al., 2017), strike a balance between these two aspects. For a comprehensive discussion of these traditional strategies, see Appendix B and recent surveys in Ren et al. (2021); Li et al. (2024a).
@@ 3.2 LLM-based Selection
With the emergence of LLMs, active learning strategies are being reimagined to exploit their powerful in-context reasoning and few-shot capabilities. In LLM-based selection, the model itself plays a dual role—both as a predictor and as a selector of informative queries. For instance, ActiveLLM Bayer and Reuter (2024) leverages an LLM to assess uncertainty and diversity in a completely unsupervised manner, making it particularly suitable in few-shot and model mismatch settings. Similarly, ActivePruneAzeemi et al. (2024) applies an LLM-driven approach to prune large unlabeled pools, reducing the computational burden of traditional acquisition functions for tasks such as translation, sentiment analysis, topic classification, and summarization. In another line of work, SelectLLM (Parkar et al., 2024) prompts LLMs directly to evaluate and rank the usefulness of unlabeled examples; the ranked instances are then refined using k-NN clustering to form effective few-shot demonstrations. Recent work by Jeong et al. (2024) further demonstrates that LLMs can generate meaningful rankings of examples, which in turn can inform fine-tuning for downstream tasks.
@@ 3.3 LLM-based Generation
Beyond selection from a fixed unlabeled set, LLMs also enable the generation of entirely new examples and labels, thereby extending the AL paradigm to an effectively infinite search space. In the following, we distinguish between generation strategies that remain within the confines of an existing unlabeled pool and those that extend beyond it.
@@
Several recent works integrate traditional selection metrics with LLM capabilities to improve query selection for few-shot learning. For example, Margatina et al. (2023) employ a combination of k-NN and perplexity-based strategies to balance uncertainty and diversity, demonstrating that uncertainty sampling may become an emergent property in larger models for in-context learning. In a similar vein, Mukherjee et al. (2024a) highlight the effectiveness of experimental design techniques, such as G-optimal design, for selecting high-impact unlabeled examples. Additionally, Diao et al. (2023) harness LLMs to generate multiple answers to a given question, using the variability among these answers as a proxy for uncertainty—albeit without venturing outside the initial dataset. Meanwhile, EAGLE (Bansal and Sharma, 2023) first samples examples based on a conditional informativeness criterion and then leverages LLMs to generate in-context labels, streamlining the annotation process.
@@ 4.1 Human Annotation
Traditional human annotation involves sending selected or generated instances to annotators for labeling, which remains the most accurate approach but is often costly. Several recent works have explored active learning strategies to optimize the annotation process. ActivePrune Azeemi et al. (2024) and CAL Du et al. (2024) reduce annotation costs by actively selecting the most valuable instances for labeling. For instance, in the case of imbalanced data, enhancing the model’s performance on the minority class can not only improve overall accuracy but also mitigate biases. One effective approach to achieving this is by increasing the size of the minority class. Providing human annotators with data that includes more minority samples has been shown to be effective Lyu et al. (2022). Active-Prompt Diao et al. (2023) and AL-Principle Margatina et al. (2023) focus on guiding LLMs by selecting instances where annotators verify final answers or prompt examples before model predictions. Furthermore, Beyond-Labels Yao et al. (2023) extends traditional annotation by collecting short rationales or natural language explanations alongside labels, improving both model interpretability and performance. Additionally, APL Muldrew et al. (2024) and BAL-PM Melo et al. (2024) incorporate human preference learning by asking annotators to compare or rank multiple model outputs. However, there are still several challenges in human annotation. Human annotator variability, as differences in expertise, cognitive biases, and annotation consistency can lead to label noise and disagreements, ultimately affecting model performance. Moreover, bias and fairness considerations in active learning are also an active area of research, as human annotations can inadvertently reflect societal biases, leading to skewed model predictions.
@@ 4.3 Hybrid
5 Termination: From Criterion to LLMs
Termination criterion in active learning loop is crucial for balancing model performance improvements with annotation costs. In LLM-based AL, termination criteria must consider not only traditional factors such as annotation budget $k$, model performance gains or uncertainty reduction, but also the variable costs associated with LLMs.
@@ 5.1 Traditional Approaches
In traditional active learning, one widely adopted strategy is to stop querying when performance improvements on a validation set fall below a predefined threshold Settles (2009). Other approaches monitor the stability of the model’s predictions or the uncertainty estimates, terminating annotation once these metrics indicate that the model has sufficiently converged Tong (2001). In addition, several theoretical frameworks provide sample complexity bounds as termination criteria, ensuring that marginal returns are decreasing Zhu et al. (2003). However, such criteria assume that the cost per annotation is uniform and homogeneous, which simplifies the stopping rule.

### v2

@@ 1 Introduction
Despite the success of traditional active learning methods, the advent of Large Language Models (LLMs) with remarkable reasoning and generation capabilities creates a new paradigm of active learning. For example, as illustrated in Figure 1, instead of solely relying on a predefined metric to query data instances, such as uncertainty Wang and Shang (2014); Diao et al. (2023) or diversity Agarwal et al. (2020); Citovsky et al. (2021), LLMs can now be used to select most informative instances after reasoning or even generate entirely new instances that are better suited for efficient model training Bayer and Reuter (2024); Parkar et al. (2024); Yang et al. (2024); Bhatt et al. (2024); Zhang and Nowak (2024). Moreover, with the collected informative data instances, LLMs also enable new data annotation schemes by collaborating with a human labeler or directly simulating a human labeler Xiao et al. (2023); Kholodna et al. (2024); Wang et al. (2024), which further reduces manual annotation efforts. LLM-based selection or generation can also help reduce training costs such as for supervised fine-tuning, in contrast to the main focus of traditional AL on reducing the labeling costs.
However, in spite of the immense potential of LLMs for active learning, particularly in high-quality data acquisition and annotation for efficient model training, existing surveys primarily focus on traditional active learning techniques, necessitating an up-to-date review of how LLMs have advanced AL in recent years. In this paper, we address this gap by presenting the first comprehensive survey of LLM-based AL techniques, which introduces a unifying taxonomy centered on the two main components of active learning: Querying (selecting or generating unlabeled instances) and Annotation (assigning labels). Table 1 and Figure 2 provide an overview of the proposed taxonomy. Table 2 further provide an intuitive comparison of existing LLM-based AL methods from the aspects of taxonomy and applications. Guided by this taxonomy, our survey organizes and systematically reviews recent works across key aspects of LLM-based active learning as follows.
@@
  • Stopping5): We discuss how recent works consider LLM costs for stopping the AL loop.
@@ 2 What is LLM-based Active Learning?
We start with basic notations and objective of traditional active learning and then introduce the LLM-based active learning loop with five main steps.
Traditional Active Learning
Let $\mathcal{U}=\{\bm{\mathrm{x}}_{i}\}_{i=1}^{N}$ be a pool of $N$ unlabeled instances, where $\bm{\mathrm{x}}_{i}\in\mathcal{X}$ are feature vectors in the input space $\mathcal{X}$. Furthermore, let =$\mathcal{L}=\{(\bm{\mathrm{x}}_{i},y_{i})\}_{i=1}^{M}$ be a labeled dataset, where $y_{i}\in\mathcal{Y}$ are the corresponding labels from the label space $\mathcal{Y}$ and the size of the labeled set $|\mathcal{L}|=M$ grows as more data instances are labeled. We are also given an annotation budget $k$ that limits the number of instances that can be labeled by the annotator. We have a target model $f_{\theta}:\mathcal{X}\to\mathcal{Y}$ parameterized by $\theta$ to be iteratively trained using the labeled data $\mathcal{L}$. Note that the target model $f_{\theta}$ can be any parameterized predictive model, regardless of its architecture. The objective of traditional AL is to efficiently select and label a subset of the unlabeled data instances $\bm{\mathrm{x}}_{i}\in\mathcal{U}$ to maximize the performance of model $f_{\theta}$ before reaching the annotation budget $k$.
@@
LLM-based AL shares a similar goal. However, it is not bounded by the unlabeled set $\mathcal{U}$, but can also use an LLM $\mathcal{M}$ to generate entirely new data instances denoted as $\bm{\mathrm{x}}^{\prime}\not\in\mathcal{U}$ as well as generating corresponding labels $y^{\prime}$ by simulating a human annotator. We define LLM $\mathcal{M}$ here as a decoder, encoder, or encoder-decoder architecture trained on a corpus of hundreds of millions to trillions of tokens following Gallegos et al. (2024). We formulate in details the LLM-based AL loop as follows.
  • Initialize: For a good starting point of the active learning loop, an LLM $\mathcal{M}$ can be used to annotate an initial set of labels or generate an initial dataset $\mathcal{L}_{\text{init}}$ to warm up the target model $f_{\theta}$. This approach overcomes the cold start problem that traditional AL methods face when there is no labeled data instance available and the initial model $f_{\theta}$ does not offer sufficient information for selecting informative data instances, especially when $f_{\theta}$ is not a pre-trained model.
  • Query: With an initialized model $f_{\theta}$, the Querying3) module is implemented to acquire the most informative data instances. Extending traditional AL methods that only select from the unlabeled set $\mathcal{U}$ with certain uncertainty or diversity metrics, the LLM $\mathcal{M}$ can now be used to select instances $\bm{\mathrm{x}}_{i}\in\mathcal{U}$ either by scoring or directly choosing, augment existing examples with generated paraphrases or contrast examples, or even generate entirely new instances $\bm{\mathrm{x}}^{\prime}\not\in\mathcal{U}$.
  • Annotate: The acquired data instances $\bm{\mathrm{x}}_{i}$ or $\bm{\mathrm{x}}^{\prime}$ are then passed to the Annotation4) module to obtain corresponding labels $y_{i}$ or $y^{\prime}$. The LLM $\mathcal{M}$ can be used either in conjunction with, to augment, or even to entirely replace the human labeler. Data instances $\bm{\mathrm{x}}_{i}\in\mathcal{U}$ that are selected from the unlabeled set are then excluded from $\mathcal{U}$. All labeled instances $(\bm{\mathrm{x}}_{i},y_{i})$ or $(\bm{\mathrm{x}}^{\prime},y^{\prime})$ are then added to the labeled set $\mathcal{L}$.
  • Train: With newly annotated data instances added to the labeled dataset $\mathcal{L}$, the target model $f_{\theta}$ is trained with a step update on its parameters $\theta$. The updated model $f_{{\theta}}$ is then used to provide information for the querying module in the next iteration before a stopping criterion is met.
  • Stop: The active learning loop reaches Stopping5) when a fixed annotation budget $k$ is reached, or when some property of the model $f_{\theta}$ being optimized, such as convergence, is satisfied. With LLMs, the budget can be not only based on the cost of human annotators, but also the cost of prompting the LLM or a combination of both.
@@
5:while not Stop$(k,f_{\theta},\mathcal{M})$ do $\triangleright$ Stopping5)
@@
8:     if $\bm{\mathrm{x}}\in\mathcal{U}$ then $\mathcal{U}~{}\leftarrow~{}\mathcal{U}\setminus\{\bm{\mathrm{x}}\}$
9:     $\mathcal{L}~{}\leftarrow~{}\mathcal{L}\cup\{(\bm{\mathrm{x}},y)\}$
10:     $f_{\theta}~{}\leftarrow~{}\texttt{Train}(\mathcal{L})$
11:return ${}f_{{\theta}^{*}},~{}\mathcal{L}$
@@ 3 Query: From Selection to Generation
Active learning fundamentally seeks to maximize model performance with minimal annotation cost by carefully selecting most informative examples. Traditionally, this process has relied on uncertainty-based and diversity-based metrics Settles (2011); Wang and Shang (2014); Geifman and El-Yaniv (2017); Citovsky et al. (2021). With the advent of LLMs, however, the paradigm is evolving from merely selecting examples from a fixed unlabeled pool to also generating new, high-value queries on demand, which can be especially helpful in reducing the training cost for supervised fine-tuning.
@@ 3.1 Traditional Selection
In this section, we briefly survey some of the key existing active learning query selection strategies. To effectively select informative examples, traditional methods capture how uncertain the model is in predicting the label of the example, i.e., uncertainty, or how different the chosen example is from the already selected examples in the labeled pool, i.e., diversity. Uncertainty-based methods, such as Least Confidence (Settles, 2009, 2011; Wang and Shang, 2014), Margin Sampling (Tong and Koller, 2001; Balcan et al., 2009; Settles, 2009), and Max-Entropy (Wang and Shang, 2014; Kremer et al., 2014; Diao et al., 2023), quantify how unsure a model is about its predictions. Complementary to these, diversity-based strategies—like CoreSet (Sener and Savarese, 2017; Geifman and El-Yaniv, 2017; Citovsky et al., 2021) and CDAL (Agarwal et al., 2020)—ensure that selected examples cover varied regions of the input space. Optimal designs (Pukelsheim, 2006) are also a form of diversity based on information gain. Hybrid approaches, including BADGE (Ash et al., 2019, 2021) and BALD (Kirsch et al., 2019; Gal et al., 2017), strike a balance between these two aspects. For a comprehensive discussion of these traditional strategies, see Appendix B and recent surveys in Ren et al. (2021); Li et al. (2024a).
@@ 3.2 LLM-based Selection
With the emergence of LLMs, active learning strategies are being reimagined to exploit their powerful in-context reasoning and few-shot capabilities. In LLM-based selection, the model itself plays a dual role—both as a predictor and as a selector of informative queries. For instance, ActiveLLM Bayer and Reuter (2024) leverages an LLM to assess uncertainty and diversity in a completely unsupervised manner, making it particularly suitable in few-shot and model mismatch settings. Similarly, ActivePrune Azeemi et al. (2024) applies an LLM-driven approach to prune large unlabeled pools, reducing the computational burden of traditional acquisition functions for tasks such as translation, sentiment analysis, topic classification, and summarization. In another line of work, SelectLLM (Parkar et al., 2024) prompts LLMs directly to evaluate and rank the usefulness of unlabeled examples; the ranked instances are then refined using k-NN clustering to form effective few-shot demonstrations. Ask-LLM Sachdeva et al. (2024) also directly prompts an instruction-tuned LLM to assess the quality of a training example. Recent work by Jeong et al. (2024) further demonstrates that LLMs can generate meaningful rankings of examples, which in turn can inform fine-tuning for downstream tasks.
@@ 3.3 LLM-based Generation
Beyond selection from a fixed unlabeled set, LLMs also enable the generation of entirely new examples and labels, thereby extending the AL paradigm to an effectively infinite search space. Such extension is fundamentally different from an earlier concept Angluin (1988) that only queries memberships for hypothetical instances. In the following, we distinguish between generation strategies that remain within the confines of an existing unlabeled pool and those that extend beyond it.
@@
Several recent works integrate traditional selection metrics with LLM capabilities to improve query selection for few-shot learning. For example, Margatina et al. (2023) employ a combination of k-NN and perplexity-based strategies, demonstrating that uncertainty sampling generally underperforms in few-shot in-context learning settings with some evidence that this may change with larger models. In a similar vein, Mukherjee et al. (2024a) highlight the effectiveness of experimental design techniques, such as G-Optimal design, for selecting high-impact unlabeled examples. Additionally, Diao et al. (2023) harness LLMs to generate multiple answers to a given question, using the variability among these answers as a proxy for uncertainty—albeit without venturing outside the initial dataset. Meanwhile, EAGLE (Bansal and Sharma, 2023) first samples examples based on a conditional informativeness criterion and then leverages LLMs to generate in-context labels, streamlining the annotation process.
@@ 4.1 Human Annotation
Traditional human annotation involves sending selected or generated instances to annotators for labeling, which remains the most accurate approach but is often costly. Several recent works have explored active learning strategies to optimize the annotation process. ActivePrune Azeemi et al. (2024) and CAL Du et al. (2024) reduce annotation costs by actively selecting the most valuable instances for labeling. For instance, in the case of imbalanced data, enhancing the model’s performance on the minority class can not only improve overall accuracy but also mitigate biases. One effective approach to achieving this is by increasing the size of the minority class. Providing human annotators with data that includes more minority samples has been shown to be effective Lyu et al. (2022). Active-Prompt Diao et al. (2023) and AL-Principle Margatina et al. (2023) focus on guiding LLMs by selecting instances where annotators verify final answers or prompt examples before model predictions. Furthermore, Beyond-Labels Yao et al. (2023) extends traditional annotation by collecting short rationales or natural language explanations alongside labels, improving both model interpretability and performance. Additionally, APL Muldrew et al. (2024) and BAL-PM Melo et al. (2024) incorporate human preference learning by asking annotators to compare or rank multiple model outputs. However, there are still several challenges in human annotation. Human annotator variability, as differences in expertise, cognitive biases, and annotation consistency can lead to label noise and disagreements, ultimately affecting model performance. Moreover, bias and fairness considerations in active learning are also an active area of research, as human annotations can inadvertently reflect societal biases, leading to skewed model predictions. There are also works on active learning in RLHF and DPO, which we refer to Section C for more details.
@@ 4.3 Hybrid
5 Stopping: From Criterion to LLMs
Stopping criterion in active learning loop is crucial for balancing model performance improvements with annotation costs. In LLM-based AL, stopping criteria must consider not only traditional factors such as annotation budget $k$, model performance gains or uncertainty reduction, but also the variable costs associated with LLMs.
@@ 5.1 Traditional Approaches
In traditional active learning, one widely adopted strategy is to stop querying when performance improvements on a validation set fall below a predefined threshold Settles (2009). Other approaches monitor the stability of the model’s predictions or the uncertainty estimates, terminating annotation once these metrics indicate that the model has sufficiently converged Tong (2001). In addition, several theoretical frameworks provide sample complexity bounds as termination criteria, ensuring that marginal returns are decreasing Zhu et al. (2003); Hanneke (2007); Dasgupta et al. (2009). However, such criteria assume that the cost per annotation is uniform and homogeneous, which simplifies the stopping rule.